OpenStudy (ddobos):

8) George has an annuity that pays $1,250 at the beginning of each quarter. If the economy grows at a rate of 3.49% annually, what is the value of the annuity if he received it in a lump sum now rather than over a period of nine years?

OpenStudy (presley551):

\[\frac{ 1 }{ i }-\frac{ 1 }{ i(1+i )^{nt}}+1\]

OpenStudy (presley551):

his pay= 1250 i=0.349/4 n=4 t=9 nt-1 is 35 do the rest